OpenStudy (anonymous):

Let f(x) = 3x - 10. Find the value of f(-3). A. 21 B. -1 C. -19

OpenStudy (the_fizicx99):

replace every x term with -3 f(-3) = 3(-3) - 10 f(-3) = -9 - 10 f(-3) = -19

OpenStudy (jdoe0001):

\(\bf f({\color{red}{ x}}) = 3{\color{red}{ x}} - 10\qquad f({\color{red}{ -3}}) = 3({\color{red}{ -3}}) - 10\)